OpenStudy (anonymous):

First four terms? a1=3: an+1= an * (n+1)

OpenStudy (anonymous):

\[a _{1}= 3 : a _{n+1} = a _{n} \times (n+1) \]

OpenStudy (jim766):

a2 = a1 *(1+1) a2 = 3(2) = 6 a3 = a2*(2+1) a2 = 6(3) = 18 try the next one

OpenStudy (anonymous):

a4= a3*(3+1) a4= 18(4)=72?

OpenStudy (anonymous):

and then the next one would be 360?

OpenStudy (jim766):

looks great!

OpenStudy (anonymous):

Thank You!